I need help setting up the second part of the equation.
Zelma is eighteen years older than her son. She was three times as old as he was one year ago. How old are they now?
I have z=18+s
Thanks:)

Answer by Menjax(62)

Z= 18 + S
Z-1 = 3S
17 + S = 3S
S = 17/2
S = 8.5
Z = 26.5